OpenStudy (anonymous):

HEEELP :( is this also a possible solution for the problem-- find all values of t such that the points A(t-1, 2t-1), B(4, 1-2t) and C(-3, 3t+5) are collinear. so y=mx+b so i substituted the x and y values to have 2t-1=m(t-1)+b 1-2t=4m+b 3t+5=-3m+b but i don't seem to have any success i solving it huhuhuhuhu

OpenStudy (anonymous):

well, solve the first two. find m and b in terms of t

OpenStudy (anonymous):

@telltoamit so do you think that the system makes sense, though?

OpenStudy (anonymous):

yea, t is like a number, say 3

OpenStudy (anonymous):

so far my answers don't end up as whole numbers... i've got t= -25±sqrt(649) _______________ 6 :(

OpenStudy (anonymous):

@telltoamit ? what do you think?

OpenStudy (anonymous):

no, no u dint hv to find t u hv to find m and b in terms of t

OpenStudy (anonymous):

Let's first deal with A and B: A(t-1, 2t-1), B(4, 1-2t) You correctly stated the equation of a straight line as y=mx+c. The gradient is calculated by dividing the change in y by the change in x: m=(1-2t-2t+1)/(4-t+1)=(2-4t)/(5-t) Now let's deal with A and C: A(t-1, 2t-1), C(-3, 3t+5) The gradient between these two is: m=(3t+5-2t+1)/(-3-t+1)=(t+6)/(-t-2) So if we equate these two gradients we get: (t+6)/(-t-2)=2(1-2t)/(5-t) Multiply to get: (t+6)(5-t)=-2(1-2t)(t+2) Expand and simplify to find t. Does this help?

OpenStudy (anonymous):

 @dydlf. here what Traxter is talking about. it's abvious that the gradient between A and B equals the gradient between B and C also equals the gradient between A and C........mAB=mBC=mAC ..... m=\[\frac{ y_2-y_1 }{ x_2-x_1 }\] u'll find a quadratic equation in (t) solve for t and let us know what u get as answers

OpenStudy (anonymous):

i forgot to label my axis sorry for that

OpenStudy (anonymous):

Yes I probably should have used a diagram to explain, thanks @hubertH :)